OpenStudy (k3du53):

Using the elimination method, what is the solution of the system? Show your work. 5x+6y=17 -10x-9y=-43 So, I found the y-coord of the solution, but I'm having trouble finding the x-coord. (x,20) 10x+12y=34 -10x-9y=-43

OpenStudy (anonymous):

10x+12y=34 (Let this be Equation A) -10x-9y=-43 (Let this be Equation B) A+B: 10x+12y+(-10x-9y)=34+(-43) 10x+12y-10x-9y=34-43 12y-9y=9 3y=9 y=3

OpenStudy (k3du53):

So now I input y into the original equation and solve for x?

OpenStudy (anonymous):

yep!

OpenStudy (k3du53):

Thank you!